Hi, I really don't understand why you have to change the ln to e while solving this equation for x. Thank you.
ln4x=2.1

Answer by Natolino1983(23)

Because, both functions (lnx) and exp(x)) are injective. That means that ln(a) = ln(b) only if a = b.
Also ln(x) in the inverse fuction of exp(x), so you can always solve an expression with one function using its inverse.
you don't know the value of 4x (becuade that's the question), but you apply the inverse function on both sides and solve the linear equation that appears.
4x =exp(2,1)
x = exp(2,1)/4

x= 2,041542478
Follow me on twitter for more @natolino_